# Exergetic Port-Hamiltonian Systems Modeling Language

### v1

@@ 6.3 Reversible components
The reversible component $(I_{\text{hk}},\,\mathcal{D}_{\text{hk}})$, as already defined in 6.8, describes the coupling between the kinetic energy domain of a piston with surface area $a$ and the two hydraulic energy domains on either side of the piston. The first condition here ensures that volume is conserved. This is satisfied since
$\left[\begin{array}[]{c}0\\ 0\\ 0\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&a\\ 0&0&-a\\ -a&a&0\end{array}\right]\,\left[\begin{array}[]{c}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}-\pi_{0}}\\ {\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}-\pi_{0}}\\ 0\end{array}\right]\,.$
@@
$\begin{split}-\mathrm{P}(\mathtt{(\mathbb{R},\,volume}))\>&=\>\mathrm{P}(a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ -\mathrm{P}((\mathbb{R},\,\mathtt{volume}))\>&=\>\mathrm{P}(-a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\,.\end{split}$
Checks for the other two non-zero entries of $L$ must pass due to its skew-symmetry. ∎
The semantics of a reversible component is simply given by its Dirac structure interpreted as a relation:
@@ 6.4 Irreversible components
Irreversible components are phenomenological models of irreversible processes with local thermodynamic equilibrium. Examples include mechanical friction, fluid viscosity, electrical conduction / resistance, thermal conduction, and mass diffusion. Such models can be defined in terms of a function that maps thermodynamic forces, which drive the relaxation process, to thermodynamic fluxes, which seek to restore equilibrium. Based on the assumption that the microscopic dynamics are reversible, Onsager found that such maps must satisfy a symmetry property Onsager (1931). This motivates our definition of an Onsager structure:
@@
$\mathcal{O}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,f\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,M(e)\,e\right\}\,,$
@@
In contrast to a Dirac structure, an Onsager structure is neither power-preserving, nor is it a locally linear relation of the port variables, since destruction of exergy, or equivalently production of entropy, is described in general by a non-linear relation. Hence, $\mathcal{O}\subset\mathcal{P}_{I}$ is not a vector subbundle of $\mathcal{P}_{I}$, but merely a fiber subbundle. The symmetry of $M(e)$ corresponds to Onsager’s reciprocal relations and its non-negative definiteness implies a non-negative exergy destruction rate.
@@
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}$
Here, $\upsilon=\mathtt{p.e}$ is the velocity and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s.e}$ is the absolute temperature at which kinetic energy is dissipated into the thermal energy domain. The phenomenological parameter $d$ could also be a function of the effort variables $e$, i.e. it could depend on the velocity and the temperature, or theoretically also the base point $x$. The above can be simplified to $\mathtt{p.f}=d\,\upsilon$ and $\mathtt{s.f}=-\frac{1}{\theta}\,d\,\upsilon^{2}$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$ cancels out in the flow variables, which here are the friction force and the entropy production rate (with a minus sign since thermal energy is leaving the component). ∎
An irreversible component is defined by an Onsager structure which satisfies two extra properties ensuring thermodynamic consistency:
@@
An irreversible component $(I,\,\mathcal{O})$ is given by an interface $I$ and an Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ such that energy and all quantities of the reference environment, except for entropy, are conserved and the time-reversal invariance property does strictly not hold. The condition for conservation of energy requires that
[…]
The condition that all quantities of the reference environment, except for entropy, are conserved requires that
[…]
@@
The condition that all quantities of the reference environment, except for entropy, are conserved is trivially satisfied. The time-reversal invariance property does strictly not hold since
$\begin{split}\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,{(\mathrm{P}(\upsilon))}^{2}\,/\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\,,\end{split}$
where we used that temperature has positive parity and velocity has negative parity. ∎
@@
$\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,\alpha\,\begin{bmatrix}\frac{\theta_{2}}{\theta_{1}}&-1\\ -1&\frac{\theta_{1}}{\theta_{2}}\end{bmatrix}\,\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}\,,$
where $\theta_{1}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s_{1}.e}$ and $\theta_{2}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s_{2}.e}$ are the absolute temperatures of the two thermal domains. Then, we can check that $(I_{\text{mf}},\,\mathcal{O}_{\text{mf}})$, defines an irreversible component. Again, the phenomenological parameter $\alpha$ could for instance be a function of the temperatures. The above can be simplified to $\mathtt{s_{1}.f}=-\frac{1}{\theta_{1}}\,\alpha\,(\theta_{2}-\theta_{1})$ and $\mathtt{s_{2}.f}=-\frac{1}{\theta_{2}}\,\alpha\,(\theta_{1}-\theta_{2})$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$ cancels out in the flow variables, which are rates of entropy change due to heat transfer. ∎
The semantics of an irreversible component is simply given by its Onsager structure interpreted as a relation:
@@ 7 Composite systems
A composite system is formed by interconnecting finitely many systems, be they primitive systems or other composite systems, according to an interconnection pattern. The data defining a composite system is therefore given by an interconnection pattern and for each of its inner boxes a system with the corresponding interface.
@@
Let $\mathbf{System}$ refer to the set of all systems defined with respect to a fixed reference environment. A composite system $(I,\,\gamma,\,J)$ is defined by the following data:
@@
    A package of systems $\gamma\in\mathbf{Pkg}(\mathbf{System})$, which can be reduced to a package of interfaces $\beta\in\mathbf{Pkg}(\mathbf{Interface})$ by forgetting further data.
@@
class of system interface relation definition
@@
Table 2 recalls the different classes of components defined in Section 6. In order to write the relations that determine their semantics in a consistent form, the Cartesian unit is inserted a few times. As already shown in the last row of the table, we can state the semantics of a composite system in the same form. The idea is that we just need to compose the relation assigned to the interconnection pattern with the Cartesian product of the relations assigned to the subsystems filling its inner boxes.
@@
Let $(I,\,\gamma,\,J)$ be a composite system and let $P=(\beta,\,I,\,J)$ be its underlying interconnection pattern. Further, let $\mathcal{P}_{P}$ be the relation assigned to pattern $P$ according to 4.2. For any inner box $b\in\operatorname{dom}(\gamma)$, let $\left\llbracket{s}\right\rrbracket$ denote the relation assigned to the subsystem $s=\gamma(b)$. In case system $s$ is itself a composite system, this definition applies recursively. Then, the semantics assigned to composite system $(I,\,\gamma,\,J)$ is the relation $\mathcal{C}:\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ given by
@@
Here, $\mathcal{X}_{I_{s}}$ denotes the state space of all storage components and $\mathcal{X}_{I_{e}}$ denotes the state space of all environment components, which are subsystems of the composite system in a recursive sense. ∎
@@
For an isolated composite system the assigned relation $\mathcal{C}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow 1$ simply amounts to a subbundle of $\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}$, which should implicitly define a vector field on $\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$. Given an initial condition, one can integrate this vector field to determine the evolution of the system. We simply say that $(x,\,\dot{x})\in\mathcal{C}$, where $x=(x_{s},\,x_{e})$, gives the dynamics of the isolated system. This is generalized by the following definition:
@@
Let $s\in\mathbf{System}$ be a system with interface $I$ and let the relation $\mathcal{R}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ define its semantics, according to one of the definitions mentioned in Table 2. Further, let $(x_{s},\,x_{e})\in\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$ denote the state variables associated to all (nested) storage and environment components. If $s$ is a primitive system this may be vacuous. Moreover, let $(x,f,e)\in\mathcal{P}_{I}$ denote the port variables of its (outer) interface. Again, if the system is isolated this is vacuous. Then, the dynamics of system $s$ is determined by the implicit differential-algebraic equation
@@
For programming languages it usually holds that not everything that is syntactically correct is also semantically meaningful. There are also cases where a compiler for the EPHS language should throw an error, despite of syntactic correctness. Every junction must have either zero or one connected port belonging to a storage or environment component. Also, one cannot connect more than two power ports with a quantity such as $(\mathbb{R},\,\mathtt{displacement})$ to the same junction, since velocities don’t balance. As of now, we don’t have a comprehensive set of restrictions that should be imposed on the interconnection of systems in order to get something meaningful, but we hope that this can be worked out in the future.

### v2

@@ 6.3 Reversible components
According to the chain rule, a Dirac structure $\mathcal{D}$ conserves the quantity corresponding to some port $p_{R}\in I_{R}$ if any resulting flow variable $f_{1}$ satisfies $0=\langle\lambda(I_{1},\,p_{R})\mid f_{1}\rangle$. Inserting the expression for $f_{1}$, this condition is satisfied if $0=\langle\lambda(I_{1},\,p_{R})\mid L(x)\,e_{1}\rangle$ is satisfied for any effort variable $e_{1}$, and $0=\langle\lambda(I_{1},\,p_{R})\mid g(x)\,f_{2}\rangle$ is satisfied for any flow variable $f_{2}$, and $0=\langle\lambda(I_{1},\,p_{R})\mid C^{*}(x)\,\lambda_{c}\rangle$ is satisfied for any Lagrange multiplier $\lambda_{c}$. Taking the linear adjoint (transpose), the first condition requires that $0=\langle L^{*}(x)\,\lambda(I_{1},\,p_{R})\mid e_{1}\rangle$ for any $e_{1}$. This simplifies to the first part of the condition in the definition, since $L$ is skew-adjoint. The other two parts follow analogously. Regarding details on the time-reversal invariance condition, we refer to [46] or [3].
@@
The reversible component $(I_{\text{hkc}},\,\mathcal{D}_{\text{hkc}})$, as already defined in 6.8, describes the coupling between the kinetic energy domain of a piston with surface area $a$ and the two hydraulic energy domains on either side of the piston. The first condition here ensures that volume is conserved. It is satisfied since
$\left[\begin{array}[]{c}0\\ 0\\ 0\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&-a\\ 0&0&+a\\ +a&-a&0\end{array}\right]\,\left[\begin{array}[]{c}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\\ {\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\\ 0\end{array}\right]\,.$
@@
$\begin{split}-\mathrm{P}(\mathtt{(\mathbb{R},\,volume}))\>&=\>\mathrm{P}(-a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ -\mathrm{P}((\mathbb{R},\,\mathtt{volume}))\>&=\>\mathrm{P}(+a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\,.\end{split}$
Checks for the other two non-zero entries of $L$ pass due to skew-symmetry of $L$. ∎
The semantics of a reversible component is simply given by its Dirac structure interpreted as a relation.
@@ 6.4 Irreversible components
Irreversible components are phenomenological models of irreversible processes with local thermodynamic equilibrium. Examples include mechanical friction, fluid viscosity, electrical conduction / resistance, thermal conduction, and mass diffusion. Such models can be defined in terms of a function that maps thermodynamic forces, which drive the relaxation process, to thermodynamic fluxes, which seek to restore equilibrium. Based on the assumption that the microscopic dynamics are reversible, Onsager found that such maps must satisfy a symmetry property [15]. This motivates our definition of an Onsager structure.
@@
$\mathcal{O}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,f\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,M(e)\,e\right\}\,,$
@@
In contrast to a Dirac structure, an Onsager structure is neither power-preserving, nor is it a locally linear relation of the port variables, since destruction of exergy, or equivalently production of entropy, is usually described by a non-linear relation. Hence, $\mathcal{O}\subset\mathcal{P}_{I}$ is not a vector subbundle of $\mathcal{P}_{I}$, but merely a fiber subbundle. The symmetry of $M(e)$ corresponds to Onsager’s reciprocal relations and its non-negative definiteness implies a non-negative exergy destruction rate, i.e. $\langle e\mid f\rangle\geq 0$.
@@
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}$
Here, $\upsilon=\mathtt{p.e}$ is the velocity and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s.e}$ is the absolute temperature at which kinetic energy is dissipated into the thermal energy domain, since adding ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels the shift $-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ in the effort variable, see 6.3. The phenomenological parameter $d$ could also be a function of the effort variables $e$, i.e. it could depend on the velocity and the temperature, or theoretically also the base point $x$. The above can be simplified to $\mathtt{p.f}=d\,\upsilon$ and $\mathtt{s.f}=-\frac{1}{\theta}\,d\,\upsilon^{2}$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels out in the flow variables, which here are the friction force and the entropy production rate (with a minus sign since thermal energy is leaving the component). The exergy destruction rate is given by $\langle\mathtt{p.e}\mid\mathtt{p.f}\rangle+\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\frac{d\,\upsilon^{2}}{\theta}\geq 0$, with $\frac{d\,\upsilon^{2}}{\theta}\geq 0$ being the entropy production rate. ∎
An irreversible component is defined by an Onsager structure that satisfies two additional conditions ensuring thermodynamic consistency.
@@
An irreversible component $(I,\,\mathcal{O})$ is given by an interface $I$ and an Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ such that energy and all quantities of the reference environment, except for entropy, are conserved and the time-reversal invariance property does strictly not hold. The condition for conservation of energy requires that $\forall\,e\in\mathrm{T}^{*}\mathcal{X}_{I,p}\,\colon$
$0\>=\>M(e)\,\bigl{(}\lambda(I)+e\bigr{)}\,.$
The condition that all quantities present in the reference environment, except for entropy, are conserved requires that $\forall\,p_{R}\in I_{R}$ with $\tau_{R}(p_{R})\neq((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$, we have that $\forall\,e\in\mathrm{T}^{*}\mathcal{X}_{I,p}\,\colon$
$0\>=\>M(e)\,\lambda(I,\,p_{R})\,.$
@@
From 6.4 and 6.3 we know that efforts variables are given by the differential of the energy function and additional shifts by the constant, intensive quantities of the environment, whenever the quantity of the respective port matches with a quantity present in the envirionment. We thus have $e=e^{\prime}-\lambda(I)$, where $e^{\prime}$ is precisely the part that stems from the differential of the energy. From this and the chain rule, it follows that energy is conserved if $0=\langle e^{\prime}\mid f\rangle=\langle e+\lambda(I)\mid M(e)\,e\rangle$. Taking the linear adjoint, this results in the condition in the definition, since $M(e)$ is symmetric. Regarding the condition that all quantities that are also present in the environment, except for entropy, are conserved, follows as already shown for reversible components, following 6.9. Again, regarding the time-reversal invariance condition, we refer to [46] or [3].
@@
The condition that all quantities of the reference environment, except for entropy, are conserved is trivially satisfied. As required, the time-reversal invariance property does strictly not hold since
$\begin{split}\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,{(\mathrm{P}(\upsilon))}^{2}\,/\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\,.\end{split}$
Here, we used that temperature has positive parity and velocity has negative parity, which follows from the parities of entropy and momentum, knowing that energy has positive parity. ∎
Irreversible components, and hence EPHS in general, respect Onsager’s reciprocal relations because the symmetry of $M(e)$ in 6.12 implies equal cross-effect coefficients (see, e.g., [48]). In the terminology of Linear Irreversible Thermodynamics (LIT), thermodynamic fluxes are related to thermodynamic forces by a symmetric operator. The following example shows this in detail.
@@
$\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\alpha\,\begin{bmatrix}\frac{\theta_{2}}{\theta_{1}}&-1\\ -1&\frac{\theta_{1}}{\theta_{2}}\end{bmatrix}\,\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}\,,$
where $\theta_{1}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s_{1}.e}$ and $\theta_{2}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s_{2}.e}$ are the absolute temperatures of the two thermal energy domains. It can be easily checked that $(I_{\text{ht}},\,\mathcal{O}_{\text{ht}})$ defines an irreversible component. Again, the phenomenological parameter $\alpha$ could for instance be a function of the temperatures. The above can be simplified to $\mathtt{s_{1}.f}=-\frac{1}{\theta_{1}}\,\alpha\,(\theta_{2}-\theta_{1})$ and $\mathtt{s_{2}.f}=-\frac{1}{\theta_{2}}\,\alpha\,(\theta_{1}-\theta_{2})$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels out in the flow variables, which are rates of entropy change due to heat transfer. Since $M(e)$ is symmetric and non-negative definite, an Onsager structure can be represented in a factorized form:
$\underbrace{\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}}_{f}\>=\>\underbrace{\begin{bmatrix}+\frac{1}{\theta_{1}}\\ -\frac{1}{\theta_{3}}\end{bmatrix}}_{C(e)}\,\underbrace{\begin{bmatrix}\alpha\,\theta_{1}\,\theta_{3}\end{bmatrix}}_{D(e)}\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,\underbrace{\begin{bmatrix}+\frac{1}{\theta_{1}}&-\frac{1}{\theta_{3}}\end{bmatrix}}_{C^{*}(e)}\,\underbrace{\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}}_{e}$
The factorization has a physical interpretation in terms of LIT (see, e.g., [13]). The thermodynamic force that drives the irreversible process is given by
$X\>=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,C^{*}(e)\,e\>=\>\frac{1}{\theta_{2}}-\frac{1}{\theta_{1}}\,.$
Application of the symmetric operator $D(e)$ yields the thermodynamic flux
$J\>=\>D(e)\,\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,C^{*}(e)\,e\>=\>\alpha\,(\theta_{1}-\theta_{2})\,.$
Since $M(e)$ models a single process, $D(e)$ has rank $1$. Application of $C(e)$ yields the flow variables $f$. The net power $\langle\mathtt{s_{1}.e}\mid\mathtt{s_{1}.f}\rangle+\langle\mathtt{s_{2}.e}\mid\mathtt{s_{2}.f}\rangle=\langle e\mid f\rangle={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\frac{\alpha\,{(\theta_{1}-\theta_{2})}^{2}}{\theta_{1}\,\theta_{2}}\geq 0$ is the exergy destruction rate, and $\frac{\alpha\,{(\theta_{1}-\theta_{2})}^{2}}{\theta_{1}\,\theta_{2}}=\langle X\mid J\rangle\geq 0$ is the entropy production rate. ∎
The semantics of an irreversible component is simply given by its Onsager structure interpreted as a relation.
@@ 7 Composite systems
A composite system is formed by interconnecting a finite number of (primitive or composite) systems according to an interconnection pattern. The data defining a composite system consists of an interconnection pattern and, for each of its inner boxes, a system with the corresponding interface.
@@
Let $\mathsf{System}$ refer to the set of all systems defined with respect to a fixed reference environment. A composite system $(I,\,\gamma,\,J)$ is defined by the following data:
@@
    A directory of systems $\gamma\in\mathsf{Dtry}(\mathsf{System})$, which can naturally be reduced to a directory of interfaces $\beta\in\mathsf{Dtry}(\mathsf{Interface})$.
@@
class of systems interface relation definition
@@
Table 2 recalls the different kinds of components defined in Section 6. In order to write the relations that determine their semantics in a unified form, the Cartesian unit is inserted a few times. As shown in the last row of the table, the semantics of a composite system takes the same form. We just need to compose the relation assigned to the interconnection pattern with the Cartesian product of the relations assigned to the subsystems filling its inner boxes.
@@
Let $(I,\,\gamma,\,J)$ be a composite system and let $P=(\beta,\,I,\,J)$ be its underlying interconnection pattern. Further, let $\mathcal{P}_{P}$ be the relation assigned to pattern $P$ according to 4.2. For any filled inner box $b\in\operatorname{dom}(\gamma)$, let $\left\llbracket{s}\right\rrbracket$ denote the relation assigned to the subsystem $s=\gamma(b)$. In case system $s$ is itself a composite system, the present definition applies recursively. Then, the semantics assigned to composite system $(I,\,\gamma,\,J)$ is the relation $\mathcal{C}:\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ given by
@@
Here, $\mathcal{X}_{I_{s}}$ denotes the state space of all storage components and $\mathcal{X}_{I_{e}}$ denotes the state space of all environment components that are subsystems of the composite system in a recursive sense. ∎
@@
For an isolated composite system the assigned relation $\mathcal{C}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow 1$ simply amounts to a subbundle of $\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}$, which implicitly defines a vector field on $\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$. Given an initial condition, one can integrate this vector field to determine the evolution of the system. We simply say that $(x,\,\dot{x})\in\mathcal{C}$, where $x=(x_{s},\,x_{e})$, gives the dynamics of the isolated system. The following definition generalizes this statement.
@@
Let $s\in\mathsf{System}$ be a system with interface $I$ and let the relation $\mathcal{R}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ define its semantics, according to the definitions mentioned in Table 2. Further, let $(x_{s},\,x_{e})\in\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$ denote the state variables associated to all (nested) storage and environment components. If the system is a reversible or an irreversible component or given by an interconnection of only such components this is vacuous. Moreover, let $(x,f,e)\in\mathcal{P}_{I}$ denote the port variables of its (outer) interface. If the system is isolated this is vacuous. Then, the dynamics of system $s$ is implicitly determined by
@@
In the context of programming languages, syntactically correct code does not necessarily define a semantically valid program. Similarly, there are cases where a compiler for the EPHS language must raise an error. For example, no junction can have two or more connected ports belonging to storage or environment components. A more comprehensive discussion of the conditions required for a composite system to be well-defined is left for future work.